You are capable of pushing a 5.2 kg box hard enough to create an acceleration of 0.38 m/s2. When you push a second box with the same force, it only accelerates at 0.21 m/s2. What is the mass of the second box?
Answer #1

Force needed the push the first box is given by

F=ma =5.2*0.38 =1.976 N

Since same force is used to push the second box ,the mass of the second box is

m2=F/a2 = 1.976/0.21

m2 = 9.41 kg